The boom is used to support the 375N flower pot. Determine the tension in wires AB and AC.
1
Expert's answer
2021-02-22T07:31:10-0500

"3j[2\/7F"ABi-6/7FABj+3/7ABk] + [-2/7FACi- 6/7FACj +3/7FACk] +[-375k]=0

-12/7FAB+12/7FAC=0

18/7FAB+18/7FAC-2250=0

FAB=FAC=437.5N